A glass tube of 1.0 m length is filled with water. The water can be drained out slowly at the bottom of the tube. If a vibrating tuning fork of frequency 500 Hz is brought at the upper end of the tube and the velocity of sound is 330 m/s, then the total number of resonance obtained will be
(a) 4
(b) 3
(c) 2
(d) 1

Grade:Upto college level

3 Answers

Chunduru Vaishnav
6 years ago
λ=V/n for closed tube, suggestions of resosances respective at λ/4, 3λ/4, 5λ/4
​so total number of resosances​ obtained​will be 3
Chunduru Vaishnav
6 years ago
λ=V/n for closed tube, suggestions of resosances respective at λ/4, 3λ/4, 5λ/4
​so total number of resosances​ obtained ​will be 3
Chunduru Vaishnav
6 years ago
λ=V/n for closed tube, suggestions of resosances respective at λ/4, 3λ/4, 5λ/4
​so total number of resosances​ obtained ​will be 3.